# Towards Sustainable Universal Deepfake Detection with Frequency-Domain Masking

### v1

@@ 4.4. Comparison with State-Of-The-Art
Chai et al. (DBLP:conf/eccv/ChaiBLI20) ResNet50 98.86 72.04 68.79 92.96 55.9 92.06 60.18 65.82 52.87 68.74 67.59 70.05 87.84 84.94 88.1 74.54 76.28 75.84 77.07 75.28
@@
Natraj et al. (Nataraj2019DetectingGG) - 99.74 80.95 50.61 98.63 53.11 67.99 59.14 68.98 60.42 73.06 87.21 70.20 91.21 89.02 92.39 89.32 88.35 82.79 80.96 78.11
Zhang et al. (Zhang2019DetectingAS) CycleGAN 55.39 100.0 75.08 55.11 66.08 100.0 45.18 47.46 57.12 53.61 50.98 57.72 77.72 77.25 76.47 68.58 64.58 61.92 67.77 66.21
Ojha et al. (ojha2023fakedetect) CLIP-RN50 100.0 99.46 99.59 97.24 99.98 99.60 82.45 61.32 79.02 96.72 99.00 87.77 99.14 92.15 99.17 94.74 95.34 94.57 97.15 93.38
Tan et al. (Tan2024RethinkingTU) S-ResNet50 100.0 90.49 84.55 99.61 77.76 100 90.15 91.21 87.34 67.59 67.72 94.81 98.39 96.94 98.56 88.65 92.66 91.22 93.58 90.06
Wang et al. (wang2019cnngenerated) ResNet500.5 100.0 96.83 88.24 98.51 98.09 95.45 66.27 92.76 63.87 98.94 99.52 68.35 65.92 66.74 65.99 72.03 76.52 73.22 66.26 81.76
@@
Wang et al. (wang2019cnngenerated) + Ours ResNet500.5 100.0 92.97 90.06 98.15 97.80 87.94 73.09 89.92 74.57 96.97 98.00 70.98 77.75 72.44 78.21 78.53 82.61 79.27 78.16 85.13
@@
Gragnaniello et al. (Gragnaniello2021) ResNet50nd 100.0 83.25 94.3 99.95 91.7 99.99 91.24 92.77 73.69 98.19 97.85 78.29 87.01 86.86 87.95 88.24 92.5 89.33 91.32 90.76
Gragnaniello et al. (Gragnaniello2021) + Ours ResNet50nd 100.0 94.07 97.61 99.92 98.54 99.99 94.0 95.46 81.47 96.72 95.47 80.14 94.35 94.24 94.77 90.32 93.46 91.05 96.29 94.10
Table 3. Mean Average Precision (mAP) comparison with state-of-the-art methods across various deepfake detection benchmarks. Bold values indicate improvements of our approach over Wang et al. (wang2019cnngenerated) and Gragnaniello et al. (Gragnaniello2021) methods. S-ResNet50 denotes a shallow ResNet architecture, ResNet500.5 indicates data processing with 50% probability of Gaussian blur and JPEG compression, and ResNet50nd means the first conv layer of ResNet was not downsampled.
@@
Chai et al. (DBLP:conf/eccv/ChaiBLI20) ResNet50 96.62 69.71 66.705 87.61 56.545 86.175 57.75 65.205 52.055 61.515 61.35 67.595 83.465 80.555 83.73 70.8 72.415 71.94 73.26 71.84
@@
Natraj et al. (Nataraj2019DetectingGG) - 98.72 72.05 52.18 95.565 52.105 61.345 58.12 66.02 58.135 69.355 76.505 65.35 80.955 79.79 81.695 79.785 78.975 76.345 74.26 72.49
Zhang et al. (Zhang2019DetectingAS) CycleGAN 52.65 99.95 62.79 52.505 58.19 99.85 47.64 48.73 52.56 52.105 50.54 54.31 64.06 63.825 63.385 60.14 57.99 56.16 58.885 60.83
Ojha et al. (ojha2023fakedetect) CLIP-RN50 100.0 98.98 97.045 89.62 99.74 98.30 74.525 62.16 68.26 78.11 85.50 78.90 96.665 82.955 96.765 86.905 87.595 86.355 91.965 87.38
Tan et al. (Tan2024RethinkingTU) S-ResNet50 100.0 86.30 82.10 98.00 70.30 100 94.60 97.90 91.80 84.50 85.10 92.90 93.60 91.90 94.00 89.30 92.80 95.10 71.10 90.59
Wang et al. (wang2019cnngenerated) ResNet500.5 100.0 93.1 85.7 98.9 94.8 96.4 85.0 97.2 75.1 98.0 98.9 77.5 75.0 72.1 75.6 83.8 87.1 85.3 59.9 85.77
@@
Wang et al. (wang2019cnngenerated) + Ours ResNet500.5 100.0 92.8 91.3 98.1 97.9 86.5 72.7 89.2 75.0 97.1 98.1 72.8 80.2 74.4 80.7 80.2 84.1 80.4 80.0 85.84
@@
Gragnaniello et al. (Gragnaniello2021) ResNet50nd 100.0 87.20 95.50 99.90 94.00 100 89.20 94.70 74.30 98.90 98.50 79.80 87.30 86.70 88.20 87.90 92.20 89.30 91.80 91.33
Gragnaniello et al. (Gragnaniello2021) + Ours ResNet50nd 100.0 93.70 98.00 99.90 98.70 100 93.00 95.70 81.80 97.30 96.20 79.50 94.20 93.80 94.80 90.00 93.10 90.90 96.40 94.00
Table 4. Area Under Receiver Operating Characteristic (AUROC) comparison with state-of-the-art methods across various deepfake detection benchmarks. Bold values indicate improvements of our approach over Wang et al. (wang2019cnngenerated) and Gragnaniello et al. (Gragnaniello2021) methods. S-ResNet50 denotes a shallow ResNet architecture, ResNet500.5 indicates data processing with 50% probability of Gaussian blur and JPEG compression, and ResNet50nd means the first conv layer of ResNet was not downsampled.
As shown in Table 3 and 4, our approach (‘+ Ours’), incorporating 15% random masking across all frequency bands, consistently enhances performance when integrated with existing state-of-the-art (SOTA) methods. Specifically, the combination of our frequency-based masking technique with Wang et al.’s method (wang2019cnngenerated) results in improvements of mAP and AUROC for all backbones used. Here, S-ResNet50 refers to a shallow ResNet architecture, ResNet500.5 indicates data processing with 50% probability of Gaussian blur and JPEG compression, and ResNet50nd denotes a modified ResNet where the first convolutional layer was not downsampled. Furthermore, integrating our method with that of Gragnaniello et al. (Gragnaniello2021) yields a performance that beats SOTA for all metrics. Notably, the gains are particularly pronounced in cross-dataset evaluations (non-ProGAN datasets), suggesting improved generalization capabilities. These results demonstrate the effectiveness of frequency masking, as it enables the model to learn more generalizable features for universal deepfake detection that are potentially overlooked by existing approaches.
@@ 4.5. Combined Augmentation Strategies
To comprehensively evaluate the generalization advantage of frequency-domain masking, we compare it against combined geometric masking strategies (Figure 4). The combination of translation and rotation—a common robust training strategy—achieves 87.18% mAP and 87.14% AUROC, showing moderate improvement over individual geometric transformations but still falling short of standalone frequency masking (88.10% mAP, 88.7% AUROC). Crucially, the integration of translation with frequency masking yields exceptional performance (90.51% mAP, 90.58% AUROC), demonstrating these strategies’ complementary nature. This suggests translation disrupts spatial artifacts while frequency masking targets spectral inconsistencies, providing more comprehensive forensic analysis. However, adding rotation to this combination (Rotate+Translate+Frequency) reduces performance to 86.97% mAP and 86.84% AUROC, indicating potential interference between rotation and spectral analysis. These results confirm that carefully selected combined strategies can outperform individual approaches, with translation-frequency masking emerging as particularly effective for universal deepfake detection.
@@ 4.7. Comparison of Frequency Transforms
We investigate alternative frequency transforms to validate our Fourier-based approach (Table 6). The Discrete Cosine Transform (DCT) slightly outperforms Fourier (88.52% mAP, 89.00% AUROC vs 88.10% mAP, 88.72% AUROC), potentially due to its energy compaction properties concentrating artifacts in fewer coefficients. Wavelet transforms underperform (86.93% mAP, 86.90% AUROC), possibly because their multi-resolution analysis dilutes generative artifacts across scales. Despite DCT’s marginal gain, we retain Fourier transforms due to its established efficacy in prior forensic work (Corvi_2023_CVPR) and minimal performance difference.
@@ 4.8. Augmentation Benefits across Pruning Levels
Slim (Liu2017LearningEC) ICCV’17 Baseline (wang2019cnngenerated) 89.93 89.12 76.03 75.15 53.95 53.47 66.65 64.65
@@
LAMP (Lee2020LayeradaptiveSF) ICLR’21 Baseline (wang2019cnngenerated) 89.93 89.12 80.63 80.85 58.80 59.02 63.36 63.96
@@
GReg (Wang2020NeuralPV) ICLR’21 Baseline (wang2019cnngenerated) 89.93 89.12 80.62 80.82 58.47 58.52 64.88 64.99
@@
DepGraph (Fang2023DepGraphTA) CVPR’23 Baseline (wang2019cnngenerated) 89.93 89.12 80.79 80.99 58.83 58.91 63.90 63.90
@@
Slim (Liu2017LearningEC) ICCV’17 Baseline (wang2019cnngenerated) 85.11 82.99 79.25 80.34 47.56 48.30 59.85 64.93
@@
LAMP (Lee2020LayeradaptiveSF) ICLR’21 Baseline (wang2019cnngenerated) 85.11 82.99 82.39 82.47 71.08 72.82 64.03 65.69
@@
GReg (Wang2020NeuralPV) ICLR’21 Baseline (wang2019cnngenerated) 85.11 82.99 83.07 83.48 72.21 74.12 66.13 67.88
@@
DepGraph (Fang2023DepGraphTA) CVPR’23 Baseline (wang2019cnngenerated) 85.11 82.99 80.89 80.85 73.49 75.55 64.48 65.59
@@
We adopt several pruning strategies including Slim (Liu2017LearningEC), LAMP (Lee2020LayeradaptiveSF), GReg (Wang2020NeuralPV), and DepGraph (Fang2023DepGraphTA). The augmentation baseline refers to training without any masking or augmentation, following the protocol of Wang et al. (wang2019cnngenerated). From Table 7, we observe consistent trends across both MobileNetv2 and ResNet50. Notably, when no pruning is applied ($p=0.0$), frequency-based augmentation achieves the highest performance, validating its regularization strength during full-capacity training.
@@ 4.9. Use Case: Synthetic Dataset from Aquaculture
Method ControlNet (Zhang2023AddingCC) Stable Diffusion (Rombach2021HighResolutionIS)
@@
Tan et al. (Tan2024RethinkingTU) 54.67 43.81 50.47 35.76
Gragnaniello et al. (Gragnaniello2021) 63.16 56.86 61.17 57.31
Wang et al. (wang2019cnngenerated) 73.38 71.46 68.69 68.77
@@
Table 8. mAP and AUROC results on the FakeFish dataset generated using ControlNet (Zhang2023AddingCC) and Stable Diffusion (Rombach2021HighResolutionIS).
@@
To explore this idea, we collected a synthetic dataset called FakeFish, generated using two popular image synthesis methods: ControlNet (Zhang2023AddingCC) and Stable Diffusion (Rombach2021HighResolutionIS). Table 8 shows our model outperforming prior work by a large margin in both mean Average Precision (mAP) and Area Under the ROC Curve (AUROC). This suggests that our approach is effective at identifying low-quality or manipulated synthetic images in aquaculture scenarios.
@@
Figure 5. Visualization of Real and AI-generated fish images from ControlNet (Zhang2023AddingCC) and Stable Diffusion (Rombach2021HighResolutionIS) in the FakeFish dataset.

### v2

@@ 4.4. Comparison with State-Of-The-Art
Chai et al. (Chai et al., 2020) ResNet50 98.86 72.04 68.79 92.96 55.9 92.06 60.18 65.82 52.87 68.74 67.59 70.05 87.84 84.94 88.1 74.54 76.28 75.84 77.07 75.28
@@
Natraj et al. (Nataraj et al., 2019) - 99.74 80.95 50.61 98.63 53.11 67.99 59.14 68.98 60.42 73.06 87.21 70.20 91.21 89.02 92.39 89.32 88.35 82.79 80.96 78.11
Zhang et al. (Zhang et al., 2019) CycleGAN 55.39 100.0 75.08 55.11 66.08 100.0 45.18 47.46 57.12 53.61 50.98 57.72 77.72 77.25 76.47 68.58 64.58 61.92 67.77 66.21
Ojha et al. (Ojha et al., 2023) CLIP-RN50 100.0 99.46 99.59 97.24 99.98 99.60 82.45 61.32 79.02 96.72 99.00 87.77 99.14 92.15 99.17 94.74 95.34 94.57 97.15 93.38
Tan et al. (Tan et al., 2024a) S-ResNet50 100.0 90.49 84.55 99.61 77.76 100 90.15 91.21 87.34 67.59 67.72 94.81 98.39 96.94 98.56 88.65 92.66 91.22 93.58 90.06
Wang et al. (Wang et al., 2020b) ResNet500.5 100.0 96.83 88.24 98.51 98.09 95.45 66.27 92.76 63.87 98.94 99.52 68.35 65.92 66.74 65.99 72.03 76.52 73.22 66.26 81.76
@@
Wang et al. (Wang et al., 2020b) + Ours ResNet500.5 100.0 92.97 90.06 98.15 97.80 87.94 73.09 89.92 74.57 96.97 98.00 70.98 77.75 72.44 78.21 78.53 82.61 79.27 78.16 85.13
@@
Gragnaniello et al. (Gragnaniello et al., 2021) ResNet50nd 100.0 83.25 94.3 99.95 91.7 99.99 91.24 92.77 73.69 98.19 97.85 78.29 87.01 86.86 87.95 88.24 92.5 89.33 91.32 90.76
Gragnaniello et al. (Gragnaniello et al., 2021) + Ours ResNet50nd 100.0 94.07 97.61 99.92 98.54 99.99 94.0 95.46 81.47 96.72 95.47 80.14 94.35 94.24 94.77 90.32 93.46 91.05 96.29 94.10
Table 3. Mean Average Precision (mAP) comparison with state-of-the-art methods across various deepfake detection benchmarks. Bold values indicate improvements of our approach over Wang et al. (Wang et al., 2020b) and Gragnaniello et al. (Gragnaniello et al., 2021) methods. S-ResNet50 denotes a shallow ResNet architecture, ResNet500.5 indicates data processing with 50% probability of Gaussian blur and JPEG compression, and ResNet50nd means the first conv layer of ResNet was not downsampled.
@@
Chai et al. (Chai et al., 2020) ResNet50 96.62 69.71 66.705 87.61 56.545 86.175 57.75 65.205 52.055 61.515 61.35 67.595 83.465 80.555 83.73 70.8 72.415 71.94 73.26 71.84
@@
Natraj et al. (Nataraj et al., 2019) - 98.72 72.05 52.18 95.565 52.105 61.345 58.12 66.02 58.135 69.355 76.505 65.35 80.955 79.79 81.695 79.785 78.975 76.345 74.26 72.49
Zhang et al. (Zhang et al., 2019) CycleGAN 52.65 99.95 62.79 52.505 58.19 99.85 47.64 48.73 52.56 52.105 50.54 54.31 64.06 63.825 63.385 60.14 57.99 56.16 58.885 60.83
Ojha et al. (Ojha et al., 2023) CLIP-RN50 100.0 98.98 97.045 89.62 99.74 98.30 74.525 62.16 68.26 78.11 85.50 78.90 96.665 82.955 96.765 86.905 87.595 86.355 91.965 87.38
Tan et al. (Tan et al., 2024a) S-ResNet50 100.0 86.30 82.10 98.00 70.30 100 94.60 97.90 91.80 84.50 85.10 92.90 93.60 91.90 94.00 89.30 92.80 95.10 71.10 90.59
Wang et al. (Wang et al., 2020b) ResNet500.5 100.0 93.1 85.7 98.9 94.8 96.4 85.0 97.2 75.1 98.0 98.9 77.5 75.0 72.1 75.6 83.8 87.1 85.3 59.9 85.77
@@
Wang et al. (Wang et al., 2020b) + Ours ResNet500.5 100.0 92.8 91.3 98.1 97.9 86.5 72.7 89.2 75.0 97.1 98.1 72.8 80.2 74.4 80.7 80.2 84.1 80.4 80.0 85.84
@@
Gragnaniello et al. (Gragnaniello et al., 2021) ResNet50nd 100.0 87.20 95.50 99.90 94.00 100 89.20 94.70 74.30 98.90 98.50 79.80 87.30 86.70 88.20 87.90 92.20 89.30 91.80 91.33
Gragnaniello et al. (Gragnaniello et al., 2021) + Ours ResNet50nd 100.0 93.70 98.00 99.90 98.70 100 93.00 95.70 81.80 97.30 96.20 79.50 94.20 93.80 94.80 90.00 93.10 90.90 96.40 94.00
Table 4. Area Under Receiver Operating Characteristic (AUROC) comparison with state-of-the-art methods across various deepfake detection benchmarks. Bold values indicate improvements of our approach over Wang et al. (Wang et al., 2020b) and Gragnaniello et al. (Gragnaniello et al., 2021) methods. S-ResNet50 denotes a shallow ResNet architecture, ResNet500.5 indicates data processing with 50% probability of Gaussian blur and JPEG compression, and ResNet50nd means the first conv layer of ResNet was not downsampled.
As shown in Table 3 and 4, our approach (‘+ Ours’), incorporating 15% random masking across all frequency bands, consistently enhances performance when integrated with existing state-of-the-art (SOTA) methods. Specifically, the combination of our frequency-based masking technique with Wang et al.’s method (Wang et al., 2020b) results in improvements of mAP and AUROC for all backbones used. Here, S-ResNet50 refers to a shallow ResNet architecture, ResNet500.5 indicates data processing with 50% probability of Gaussian blur and JPEG compression, and ResNet50nd denotes a modified ResNet where the first convolutional layer was not downsampled. Furthermore, integrating our method with that of Gragnaniello et al. (Gragnaniello et al., 2021) yields a performance that beats SOTA for all metrics. Notably, the gains are particularly pronounced in cross-dataset evaluations (non-ProGAN datasets), suggesting improved generalization capabilities. These results demonstrate the effectiveness of frequency masking, as it enables the model to learn more generalizable features for universal deepfake detection that are potentially overlooked by existing approaches.
@@ 4.5. Combined Augmentation Strategies
To comprehensively evaluate the generalization advantage of frequency-domain masking, we compare it against combined geometric masking strategies (Figure 4). The combination of translation and rotation—a common robust training strategy—achieves 87.18% mAP and 87.14% AUROC, showing moderate improvement over individual geometric transformations but still falling short of standalone frequency masking (88.10% mAP, 88.7% AUROC). We believe this is because rotation is a stronger and more disruptive operation than translation: rotating faces changes their orientation and introduces interpolation artifacts, which may partly “wash out” the subtle traces that the detector needs to learn. In contrast, translation mostly shifts the content without heavily distorting local structures, which seems more helpful for robust training.
Crucially, the integration of translation with frequency masking yields exceptional performance (90.51% mAP, 90.58% AUROC), suggesting that these two augmentations act in a complementary way: translation mainly breaks simple spatial patterns, while frequency masking targets inconsistencies in the spectral domain, so together they provide a richer set of examples. However, adding rotation to this combination (Rotate+Translate+Frequency) reduces performance to 86.97% mAP and 86.84% AUROC. A simple explanation is that stacking rotation on top of translation and frequency masking may over-perturb the images, making them less similar to realistic test data and encouraging the model to rely on artificial artifacts created by the augmentations rather than on genuine deepfake cues. Overall, these results confirm that carefully selected combined strategies can outperform individual approaches, with translation–frequency masking emerging as particularly effective for universal deepfake detection.
@@ 4.7. Comparison of Frequency Transforms
We investigate alternative frequency transforms to validate our Fourier-based approach (Table 6). The Discrete Cosine Transform (DCT) slightly outperforms Fourier (88.52% mAP, 89.00% AUROC vs 88.10% mAP, 88.72% AUROC), potentially due to its energy compaction properties concentrating artifacts in fewer coefficients. Wavelet transforms underperform (86.93% mAP, 86.90% AUROC), possibly because their multi-resolution analysis dilutes generative artifacts across scales. Despite DCT’s marginal gain, we retain Fourier transforms due to its established efficacy in prior forensic work (Corvi et al., 2023a) and minimal performance difference.
@@ 4.8. Augmentation Benefits across Pruning Levels
Slim (Liu et al., 2017) ICCV’17 Baseline (Wang et al., 2020b) 89.93 89.12 76.03 75.15 53.95 53.47 66.65 64.65
@@
LAMP (Lee et al., 2020) ICLR’21 Baseline (Wang et al., 2020b) 89.93 89.12 80.63 80.85 58.80 59.02 63.36 63.96
@@
GReg (Wang et al., 2020a) ICLR’21 Baseline (Wang et al., 2020b) 89.93 89.12 80.62 80.82 58.47 58.52 64.88 64.99
@@
DepGraph (Fang et al., 2023) CVPR’23 Baseline (Wang et al., 2020b) 89.93 89.12 80.79 80.99 58.83 58.91 63.90 63.90
@@
Slim (Liu et al., 2017) ICCV’17 Baseline (Wang et al., 2020b) 85.11 82.99 79.25 80.34 47.56 48.30 59.85 64.93
@@
LAMP (Lee et al., 2020) ICLR’21 Baseline (Wang et al., 2020b) 85.11 82.99 82.39 82.47 71.08 72.82 64.03 65.69
@@
GReg (Wang et al., 2020a) ICLR’21 Baseline (Wang et al., 2020b) 85.11 82.99 83.07 83.48 72.21 74.12 66.13 67.88
@@
DepGraph (Fang et al., 2023) CVPR’23 Baseline (Wang et al., 2020b) 85.11 82.99 80.89 80.85 73.49 75.55 64.48 65.59
@@
We adopt several pruning strategies including Slim (Liu et al., 2017), LAMP (Lee et al., 2020), GReg (Wang et al., 2020a), and DepGraph (Fang et al., 2023). The augmentation baseline refers to training without any masking or augmentation, following the protocol of Wang et al. (Wang et al., 2020b). From Table 7, we observe consistent trends across both MobileNetv2 and ResNet50. Notably, when no pruning is applied ($p=0.0$), frequency-based augmentation achieves the highest performance, validating its regularization strength during full-capacity training.
@@ 4.9. Use Case: Synthetic Dataset from Aquaculture
Method ControlNet (Zhang et al., 2023) Stable Diffusion (Rombach et al., 2021b)
@@
Tan et al. (Tan et al., 2024a) 54.67 43.81 50.47 35.76
Gragnaniello et al. (Gragnaniello et al., 2021) 63.16 56.86 61.17 57.31
Wang et al. (Wang et al., 2020b) 73.38 71.46 68.69 68.77
@@
Table 8. mAP and AUROC results on the FakeFish dataset generated using ControlNet (Zhang et al., 2023) and Stable Diffusion (Rombach et al., 2021b).
@@
To explore this idea, we collected a synthetic dataset called FakeFish, generated using two popular image synthesis methods: ControlNet (Zhang et al., 2023) and Stable Diffusion (Rombach et al., 2021b). Table 8 shows our model outperforming prior work by a large margin in both mean Average Precision (mAP) and Area Under the ROC Curve (AUROC). This suggests that our approach is effective at identifying low-quality or manipulated synthetic images in aquaculture scenarios.
@@
Figure 5. Visualization of Real and AI-generated fish images from ControlNet (Zhang et al., 2023) and Stable Diffusion (Rombach et al., 2021b) in the FakeFish dataset.